OpenStudy (anonymous):

say you drop a ball 25 meters above the ground the ball bounces 75% of the previous height with each bounce. Find the total number of metres the ball travels between the time it is dropped and the 10th bounce

OpenStudy (anonymous):

leave the first 25 alone, add that in at the end goes up \(.75\times25\)and down \(.75\times 25\)for a total of \(50\times .75\) then up and down \(50\times .75^2\) etc add \[50\times .75+50\times .75^2+50\times 275^3+...+50\times .75^9\]